# CryoFM: A Flow-based Foundation Model for Cryo-EM Densities

### v1

@@ E.2 Ab initio Modeling
a pretrained vector field $v_{\theta}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, $K$ obseravations ${\mathbf{y}}^{(k)}\in\mathbb{R}^{m}$, number of steps $N$, maximum step size $\lambda_{\text{max}}$ at each step, the loss function with respect to the likelihood term $\mathcal{L}$
@@
for $t\in[1,1-\frac{1}{N},1-\frac{2}{N},\cdots,0]$ do
     ${\mathbf{x}}^{\prime}_{t-1}\leftarrow{\mathbf{x}}_{t}-v_{\theta}(t,{\mathbf{x}}_{t})$
@@
     ${\mathbf{x}}_{t-1}\leftarrow{\mathbf{x}}^{\prime}_{t-1}+\lambda_{t}{\mathbf{g}}$
@@
Appendix F Additional Ablation Study
F.1 Data Augmentation
@@
F.2 Patchifying & Downsampling Factor
@@
In Fig. 15, we present the loss curves on the test set during training for different model variants. Given the substantial differences in the computational costs among these model variants, the horizontal axis represents the training compute in Gflops. The estimation of training compute is: model Gflops $\cdot$ batch size $\cdot$ training steps $\cdot$ 3 (Peebles & Xie, 2023a). We observe that the models with higher computational costs achieve lower loss levels, indicating the potential for scalability. On the other hand, as shown in Tab. 7, the inference speeds of the p2 and d0 models are exceedingly slow. Therefore, we selected the p4-d1 configuration with a comprehensive consideration of model performance and efficiency in downstream applications.
@@
F.3 Hyperparameters for Posterior Sampling
Posterior sampling process involves continuously moving forward and correcting in the directions of prior $p({\mathbf{x}})$ and the likelihood $p({\mathbf{x}}|{\mathbf{y}})$. In this section, we present ablation experiments on the two parameters that significantly impact the effectiveness of the posterior sampling algorithm: the number of total timesteps, and $\lambda_{\text{max}}$ which controls the extent to which $x_{t}$ is modified based on the observed data.
@@ F.3 Hyperparameters for Posterior Sampling
F.4 Latent Flow Models
In this section, we discuss the choice of the input space of the flow-matching model: should we learn in the voxel space or the latent space? Most of the early works (Goodfellow et al., 2014; Van Den Oord et al., 2016; Brock, 2018; Dhariwal & Nichol, 2021) for generative modeling learned the data distribution directly on the original image space. Rombach et al. (2022) pointed out that learning diffusion models on the latent space can generate high-quality images with much less computational cost. The key is to learn an variational auto-encoder (VAE) (Kingma, 2014) which provides a low-dimensional representation that is equivalent to the original data space (Rombach et al., 2022; Podell et al., 2023). Then, a Diffusion Transformer (DiT) (Peebles & Xie, 2023a) is trained on the latent codes of the VAE.
@@ F.4.1 Architecture of LFM
  • a DiT (Diffusion Transformer) (Peebles & Xie, 2023a) that learns the distribution of the latent codes. We experimented with a standard DiT architecture by replacing the 2D patchifying module with a 3D module  [^8] https://github.com/huggingface/diffusers. The DiT has $12$ layers, each of which has $1024$ dimensions, and the patchify factor is set to $2$.
Section F.4.2 shows that VAE-$8$ performs significantly worse in reconstrution tasks since it is over-compressed and causes loss of information. However, a larger downsampling factor produces smaller number of latent codes, which may ease the training of DiT. Therefore, we trained two DiTs on the latent codes of two VAEs. For each setting, we trained the DiT for $300$k steps.
F.4.2 Evaluation of VAE

### v2

@@ G.2 Ab initio Modeling
a pretrained vector field ${\textnormal{v}}_{\Theta}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, $K$ obseravations ${\mathbf{y}}^{(k)}\in\mathbb{R}^{m}$, number of steps $N$, maximum step size $\lambda_{\text{max}}$ at each step, the loss function with respect to the likelihood term $\mathcal{L}$
@@
for $t\in[1,1-\Delta t,1-2\Delta t,\cdots,\Delta t]$ do
     ${\mathbf{x}}^{\prime}_{t-\Delta t}\leftarrow{\mathbf{x}}_{t}-\Delta t\cdot{\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})$
@@
     ${\mathbf{x}}_{t-\Delta t}\leftarrow{\mathbf{x}}^{\prime}_{t-\Delta t}-\lambda_{t}{\mathbf{g}}$
@@
Appendix H Additional Ablation Study
H.1 Data Augmentation
@@
H.2 Patchifying & Downsampling Factor
@@
In Fig. 17, we present the loss curves on the test set during training for different model variants. Given the substantial differences in the computational costs among these model variants, the horizontal axis represents the training compute in Gflops. The estimation of training compute is: model Gflops $\cdot$ batch size $\cdot$ training steps $\cdot$ 3 (Peebles & Xie, 2023). We observe that the models with higher computational costs achieve lower loss levels, indicating the potential for scalability. On the other hand, as shown in Tab. 7, the inference speeds of the p2 and d0 models are exceedingly slow. Therefore, we selected the p4-d1 configuration with a comprehensive consideration of model performance and efficiency in downstream applications.
@@
H.3 Hyperparameters for Posterior Sampling
Posterior sampling process involves continuously moving forward and correcting in the directions of prior $p({\mathbf{x}})$ and the likelihood $p({\mathbf{x}}|{\mathbf{y}})$. In this section, we present ablation experiments on the two parameters that significantly impact the effectiveness of the posterior sampling algorithm: the number of total timesteps, and $\lambda_{\text{max}}$ which controls the extent to which ${\mathbf{x}}_{t}$ is modified based on the observed data. We evaluated the performance using only the first four samples from the test set.
@@ H.3 Hyperparameters for Posterior Sampling
H.4 Latent Flow Models
In this section, we discuss the choice of the input space of the flow-matching model: should we learn in the voxel space or the latent space? Most of the early works (Goodfellow et al., 2014; Van Den Oord et al., 2016; Brock, 2018; Dhariwal & Nichol, 2021) for generative modeling learned the data distribution directly on the original image space. Rombach et al. (2022) pointed out that learning diffusion models on the latent space can generate high-quality images with much less computational cost. The key is to learn an variational auto-encoder (VAE) (Kingma, 2014) which provides a low-dimensional representation that is equivalent to the original data space (Rombach et al., 2022; Podell et al., 2023). Then, a Diffusion Transformer (DiT) (Peebles & Xie, 2023) is trained on the latent codes of the VAE.
@@ H.4.1 Architecture of LFM
  • a DiT (Diffusion Transformer) (Peebles & Xie, 2023) that learns the distribution of the latent codes. We experimented with a standard DiT architecture by replacing the 2D patchifying module with a 3D module  [^6] https://github.com/huggingface/diffusers. The DiT has $12$ layers, each of which has $1024$ dimensions, and the patchify factor is set to $2$.
Section H.4.2 shows that VAE-$8$ performs significantly worse in reconstrution tasks since it is over-compressed and causes loss of information. However, a larger downsampling factor produces smaller number of latent codes, which may ease the training of DiT. Therefore, we trained two DiTs on the latent codes of two VAEs. For each setting, we trained the DiT for $300$k steps.
H.4.2 Evaluation of VAE